OpenStudy (anonymous):

evaluate the expression for a=15,b=25,c=26,d=1 a+(b+c) x d

OpenStudy (anonymous):

plug in the values into your expression

OpenStudy (anonymous):

so you would start with a 15+(b+c) xd

OpenStudy (anonymous):

x d***

OpenStudy (anonymous):

so wut would it be?

OpenStudy (anonymous):

You tell me

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

Keep subsituting the letters for the numbers they are equal to

OpenStudy (anonymous):

its 66 if you just want the answer

OpenStudy (anonymous):

evaluate the expression for a=11,b=8,c=2,d=7 a+(b+c) x d

OpenStudy (anonymous):

What do you think the answer should be?

OpenStudy (anonymous):

if i knew i wouldnt be posting the question up

OpenStudy (anonymous):

Well do you care how to solve it or do you just want the answer. DO you have any idea on what to do? Is there somthing you think might work?

OpenStudy (slaaibak):

Simple example: If a = 15 what is 2a? Use this logic to solve the problem

OpenStudy (across):

Hello kids! Today we're going to learn how to change the names of things! Yay! Let's look at this extraneous specimen:\[a=11,b=8,c=2,d=7,\]\[a+(b+c)\cdot d.\] :O !!! Pikachu! Substitute 'a' NOW!!!\[11+(b+c)\cdot d.\] >:D Well done Pikachu, return! Bulbasaur, substitute 'b' NOW!!!\[11+(8+c)\cdot d\] ... and so on...

OpenStudy (slaaibak):

So much win!

OpenStudy (anonymous):

@ across do u have facebook?

OpenStudy (anonymous):

AWESOME!

OpenStudy (anonymous):

wuts ur email? or ur name on facebook?

OpenStudy (anonymous):

subistution is my favorite thing in math now

OpenStudy (across):

hahaha

OpenStudy (anonymous):

ahaha ill try

OpenStudy (anonymous):

how about u add me :)

OpenStudy (slaaibak):

Pick up on a math forum. NICE!

OpenStudy (anonymous):

dude i dont know how to do the math man

OpenStudy (slaaibak):

ah ye sorry. facebook will be the best solution then :)

OpenStudy (anonymous):

ahahahahaha